A rectangles length and width are in a ratio of 6:1. The area is 150 square feet. What are the length and width?
Answer by lwsshak3(11628)

A rectangles length and width are in a ratio of 6:1. The area is 150 square feet. What are the length and width?
***
let x=width of rectangle
let y=length of rectangle
y/x=6/1
y=6x
length*width=area
y*x=150
6x*x=150
6x^2=150
x^2=150/6=25
x=5
y=6*5=30
width of rectangle=5 ft
length of rectangle=30 ft